in a certain fraction, the denominator is 2 less than the numerator. if 1 is added to both the numerator and the denominator the resulting fraction will equal to 5 over 4. find the original fraction
Answer by unlockmath(1688)

Hello,
Let's use x as the number and set it up like this:
(x+1)/[(x-2)+1]=5/4
Simplify this:
(x+1)/(x-1)=5/4
Multiply by 4(x-1)to get:
4(x+1)=5(x-1)
Expand to:
4x+4=5x-5
Subtract 4x and add 5:
x=9
So now we know the fraction is:
9/7
Make sense?
RJ
www.math-unlock.com